# Meta-Computing Enhanced Federated Learning in IIoT: Satisfaction-Aware Incentive Scheme via DRL-Based Stackelberg Game

### v2

@@ V-B Stackelberg Game Analysis
We analyze the node’s optimal decision using a typical inverse induction method. Subsequently, we compute the first-order and second-order derivatives of ${U_{i}}$ with respect to ${\theta_{i}}$ as follows:
@@
Bringing $\theta_{i}^{*}$ into Eq. (10), $V$ re-expresses as
@@
where $a>1$, ${{\theta}_{i}}>at$, $\theta_{i}^{*}=\frac{{{\sigma}_{i}}}{{{r}_{i}}}>at$, and then $r<\frac{{{\sigma}_{i}}}{at}<\frac{{{\sigma}_{i}}}{(a-1)t}$, we obtain $\Lambda=\sigma-\left(a-1\right)tr>0$. The second order derivative $\frac{{{\partial}^{2}}V_{i}}{\partial{{r}_{i}}^{2}}<0$ is finally obtained. It shows that the objective function in the problem is convex with respect to ${{r}_{i}}$ and the problem in Eq. (17) is a convex optimization problem, which can be solved by existing convex optimization tools to find the optimal solution of ${{r}_{i}^{*}}$
@@ VI-A DRL for Stackelberg Game
Traditional heuristic algorithms require full information about the game environment. However, due to the non-cooperative relationship, each game player is not willing to disclose its private information. DRL aims to learn decision-making based on past experiences, current states, and given rewards. To address decision-making problems with continuous action spaces, this paper employs the MADDPG algorithm [34], a DRL algorithm for use in a multi-intelligence environment. We describe the detailed definition of each term as follows.
@@ VI-B MADDPG Algorithm
The process of the MADDPG-based Stackelberg game is illustrated in Fig. 2, where the server serves as the leader and nodes act as the followers. In each training cycle, the server agent observes the state $o^{t}_{server}$ and determines the action $r^{t}_{server}$, while the node agent observes the state $o^{t}_{node}$ and determines the action $\theta^{t}_{node}$. Subsequently, the current state transitions to the next, and the agents receives the reward. The detailed components of the DRL controller for each agent are shown in Fig. 3. The replay buffer is used to store transitions, including the current state, action, reward, and next state, collected during interactions with the environment. These stored transitions are sampled in batches to decorrelate sequential data and stabilize the training process. The actor and critic network comprising three fully connected layers. The actor network takes the current state as input and outputs the corresponding action by generating a policy. The critic network evaluates the action taken by the actor network and provides a value estimation to guide policy improvement. Both the actor and critic networks are updated using two separate optimization modules. The policy optimizer updates the parameters of the actor network based on the policy gradient, while the value optimizer minimizes the temporal difference error to refine the critic network’s value estimations.
@@
6:     Initialize a random process $N$ for action exploration
@@
12:        Sample a random minibatch of $N$ transitions $(o_{i},x_{i},e_{i},o_{i+1})$ from $B$
@@
$\begin{split}\nabla_{\phi^{\mu}}J\approx E_{o_{t}\sim\rho^{\beta}}\left[\nabla_{\phi^{\mu}}Q\left(o_{t},\mu\left(o_{t}\middle|\phi^{\mu}\right)\middle|\phi^{Q}\right)\right],\end{split}$ (20)
@@
$\begin{split}L\left(\phi^{Q}\right)=E_{\left(o_{t},x_{t},e_{t},o_{t+\mathbb{1}}\right)\sim\mathbb{D}}\left[\left(y_{t}-Q\left(o_{t},x_{t}\middle|\phi^{Q}\right)\right)^{2}\right],\end{split}$ (21)
@@
Total number of nodes $I$ [5,25]
@@ VII Simulation Results
To verify the impact of the proposed satisfaction-aware incentive scheme on FL in the IIoT, this paper conducts experiments using Python 3.7 and TensorFlow 1.15, and evaluates the performance of the scheme. The value ranges of the simulation parameters are shown in Table II. We take image classification tasks as an example of IIoT applications. In the physical space, the robots perform image classification tasks without loss of generality. We use the MNIST dataset, consisting of 60,000 samples for the training set and 10,000 samples for the test set. To evaluate the performance of the proposed scheme, we compare it with four reinforcement learning algorithms: MAPPO [35], MASAC [36], and MADQN [37], which are DDPG algorithm, PPO algorithm, SAC algorithm, and DQN algorithm in multi-agent environment.
@@ VII-A Satisfaction and Utility Analysis
Fig. 5 shows a comparative analysis of server utility under different incentive schemes with the maximum budget constraint ($R^{max}=50$). The utility of our scheme reaches its peak when 15 nodes are chosen, indicating an optimal balance between the number of nodes and the budget constraint. While increasing the number of nodes usually results in higher utilities, due to the budget limit, utilities decrease after the optimal point. The results suggest that our scheme consistently outperforms other strategies across all node numbers, indicating an optimized balance between cost and quality incentives to maximize utility within the given budget. The quality-first and price-first strategies show competitive performance, although neither dominates across all node numbers. The randomized pricing and randomized node combination strategies yield lower utility, which may suggest a less effective node selection or pricing model under these approaches.
Fig. 6 illustrates the comparison of normalized server utility under different budget constraints. The figure shows three bar sets corresponding to different maximum budget constraints: $R^{max}=50$, $R^{max}=100$, and $R^{max}=150$. The number of selected nodes ranges from 5 to 25. As the number of selected nodes increases, the server utility generally increases under each budget constraint. However, the extent of the utility increase varies with the budget limit. For the lowest budget ($R^{max}=50$), the utility increases with the number of nodes but then levels off, indicating that beyond a certain point, additional nodes do not increase utility within the budget constraint. As the budget constraint increases from $R^{max}=100$ to $R^{max}=150$, utility continues to increase with the number of nodes, suggesting that higher budgets can effectively utilize more nodes to generate utility. The highest utility is observed under the constraint $R^{max}=150$, particularly with 100 nodes. This means that the server can utilize a higher budget to maximize its utility potential. The server’s ability to generate utility is significantly affected by the number of nodes it can support, which in turn is limited by its budget.
Fig. 7 illustrates the comparative analysis of server strategies under different conditions. The unit cost per node ranges from 1 to 5. Fig. 7 shows the server bids under different unit satisfaction margins $\sigma$. As the node unit cost increases, the server bids decrease for all values of $\sigma$. Higher satisfaction margins result in a sharp decrease in bids. This suggests that as the node cost increases, servers are inclined to bid higher for nodes with higher satisfaction margins. Fig. 7 shows the server’s bid for different data sizes $d$. As in Fig. 7, all three lines exhibit a decreasing trend, indicating that as node cost increases, the server’s willingness to bid declines. This decrease is more significant at larger data volumes, implying that servers favor nodes with larger data sizes when bidding. Fig. 7 shows that both unit satisfaction profit and data size significantly affect server strategy. As the node cost increases, servers tend to place lower bids, and higher satisfaction and larger data sizes amplify this effect.
@@
Figure 7: Normalized server bids for different parameters and unit cost of node.
@@ VII-B FL Performance Comparison
Figure 8: FL training time for different number of selected nodes.
Figure 9: FL accuracy for different number of selected nodes.
@@ VII-C Solution Comparison
Fig. 10 compares the performance of different DRL algorithms in the Stackelberg game. In Fig. 10, as the number of selected clients increases, the server utility exhibits a rise-then-fall trend, reaching its peak at 15 clients. This trend occurs because a moderate number of clients provides sufficient computational resources to optimize the training performance. However, as the number of clients continues to increase, the limited total budget results in less resource allocation per node, leading to a decline in overall utility. Of the four algorithms, MADDPG consistently achieves the highest utility, closely approximating the SE. In contrast, the utilities of MAPPO, MASAC, and MADQN are slightly lower, showing a noticeable gap. This demonstrates that MADDPG excels in balancing resource allocation and client selection to maximize server utility. In Fig. 10, as the node unit cost increases, server bids gradually decrease because the server must lower its bids to optimize resource allocation under budget constraints. The server bids in MADDPG remain consistent with the SE, whereas the server bids in MAPPO, MASAC, and MADQN diverge from the SE. Fig. 10 shows that MADDPG consistently approximates the SE more closely than MAPPO, MASAC, and MADQN, indicating its superior performance in this Stackelberg game setting.
@@ VIII Conclusion
In this paper, we have designed the MEFL framework for IIoT. The framework leverages meta-computing to enhance FL in IIoT. We have developed a satisfaction-aware incentive scheme within MEFL. This scheme integrates a satisfaction function that incorporates data size, AoI, and latency into the utility function, and formulates the utility optimization problem as a two-stage Stackelberg game. We employ a DRL-based approach to learn the equilibrium of the Stackelberg game, aiming to maximize the server’s utility by identifying the optimal equilibrium. Experimental results demonstrate that, compared to traditional methods, the scheme effectively balances model quality and update latency through efficient resource allocation, enhances FL performance, ensures real-time capability and high efficiency of FL in IIoT, and better addresses the specific needs of industrial scenarios. In future work, we will extend our meta-computing framework to incorporate efficient asynchronous FL, with the aim of further improving learning efficiency in IIoT.

### v3

@@ V-B Stackelberg Game Analysis
The node’s optimal decision is analyzed using the standard inverse induction method. Subsequently, we compute the first-order and second-order derivatives of $U_{i}$} with respect to $\theta_{i}$} as follows:
@@
By substituting $\theta_{i}^{*}$ into Eq.(10),the server’s utility $V$ can be re-expressed as
@@
where $a>1$, ${{\theta}_{i}}>at$, $\theta_{i}^{*}=\frac{{{\sigma}_{i}}}{{{r}_{i}}}>at$, and then $r<\frac{{{\sigma}_{i}}}{at}<\frac{{{\sigma}_{i}}}{(a-1)t}$, we obtain $\Lambda=\sigma-\left(a-1\right)tr>0$. The second order derivative $\frac{{{\partial}^{2}}V_{i}}{\partial{{r}_{i}}^{2}}<0$ is finally obtained. It shows that the objective function in the problem is convex with respect to ${r}_{i}$} and the problem in Eq. (17) is a convex optimization problem, which can be solved by existing convex optimization tools to find the optimal solution of ${r}_{i}^{*}$} .
@@
As each IIoT node is modeled as a rational agent seeking to maximize its own utility, it is unlikely to intentionally harm the system, since such behavior would ultimately reduce its expected reward under the Stackelberg equilibrium. Even when nodes attempt limited strategic misreporting, the server can still infer their true state through observable indicators such as AoI, service latency, and participation frequency. Consequently, the proposed scheme demonstrates robustness against unreliable nodes and mild strategic behaviors.
@@ VI-A DRL for Stackelberg Game
Traditional heuristic algorithms require full information about the game environment. However, due to the non-cooperative nature of the game, players are typically unwilling to disclose their private information. DRL aims to learn decision-making based on past experiences, current states, and given rewards. To address decision-making problems with continuous action spaces, this paper employs the multi-agent deep deterministic policy gradient (MADDPG) algorithm [b55], a DRL algorithm designed for multi-agent environments. We describe the detailed definition of each term as follows.
@@ VI-B MADDPG Algorithm
The process of the MADDPG-based Stackelberg game is illustrated in Fig. 2, where the server serves as the leader and nodes act as the followers. In each training cycle, the server agent observes the state $o^{t}_{server}$ and determines the action $x^{t}_{server}$, while the node agent observes the state $o^{t}_{node}$ and determines the action $\theta^{t}_{node}$. Subsequently, the current state transitions to the next, and the agents receives the reward. The detailed components of the DRL controller for each agent are shown in Fig. 3. The replay buffer is used to store transitions, including the current state, action, reward, and next state, collected during interactions with the environment. These stored transitions are sampled in batches to decorrelate sequential data and stabilize the training process. The actor and critic network comprising three fully connected layers. The actor network takes the current state as input and outputs the corresponding action by generating a policy. The critic network evaluates the action taken by the actor network and provides a value estimation to guide policy improvement. Both the actor and critic networks are updated using two separate optimization modules. The policy optimizer updates the parameters of the actor network based on the policy gradient, while the value optimizer minimizes the temporal difference error to refine the critic network’s value estimations.
@@
6:  Initialize a random process NN for action exploration
@@
12:   Sample a random minibatch of NN transitions $(o_{i},x_{i},e_{i},o_{i+1})$ from $B$
@@
$\begin{split}\nabla_{\phi^{\mu}}J\approx\mathbb{E}_{o_{t}\sim\rho^{\beta}}\left[\nabla_{\phi^{\mu}}Q\left(o_{t},\mu\left(o_{t}\middle|\phi^{\mu}\right)\middle|\phi^{Q}\right)\right],\end{split}$ (20)
@@
$\begin{split}L\left(\phi^{Q}\right)=\mathbb{E}_{\left(o_{t},x_{t},e_{t},o_{t+1}\right)\sim\mathbb{D}}\left[\left(y_{t}-Q\left(o_{t},x_{t}\middle|\phi^{Q}\right)\right)^{2}\right],\end{split}$ (21)
@@
In addition, as the number of participating nodes $I$ increases, the parallel structure of the actor–critic networks allow each agent to update its policy independently, while the centralized critic aggregates global feedback with a linear communication cost $O(I)$. This decentralized training and centralized evaluation architecture ensures that the system’s overall computational growth remains approximately linear rather than exponential. Regarding time varying network conditions, the continuous online update of the server’s strategy and each node’s update cycle enables adaptive re-optimization without retraining from scratch. The DRL agents learn dynamic mappings from observed states to actions, allowing the system to automatically adjust to variations in communication delays, node failures, or changing participation frequencies. Consequently, the proposed method maintains convergence stability and utility optimization even when applied to real world IIoT deployments with hundreds of heterogeneous nodes and dynamically evolving conditions.
@@
Total number of nodes II [5,25]
@@ VII Simulation Results
To verify the impact of the proposed satisfaction-aware incentive scheme on FL in the IIoT, this paper conducts experiments using Python 3.7 and TensorFlow 1.15, and evaluates the performance of the scheme. The value ranges of the simulation parameters are shown in Table II. We take image classification tasks as an example of IIoT applications. In the physical space, the robots perform image classification tasks without loss of generality. We use the MNIST dataset, consisting of 60,000 samples for the training set and 10,000 samples for the test set. To evaluate the performance of the proposed scheme, we compare it with four reinforcement learning algorithms: multi-agent proximal policy optimization (MAPPO) [b53], multi-agent soft actor-critic (MASAC) [b54], and multi-agent deep Q-networks (MADQN) [b56], which are PPO algorithm, SAC algorithm, and DQN algorithm in multi-agent environment.
@@ VII-A Satisfaction and Utility Analysis
Fig. 5 shows a comparative analysis of server utility under different incentive schemes with the maximum budget constraint ($R^{max}=50$). The utility of our scheme reaches its peak when 15 nodes are chosen, indicating an optimal balance between the number of nodes and the budget constraint. While increasing the number of nodes usually results in higher utilities, due to the budget limit, utilities decrease after the optimal point. However, this does not mean that our scheme cannot handle larger scale networks. The peak point corresponds to the equilibrium where the marginal utility of adding new participants is balanced by the fixed incentive budget. As more nodes join, the average reward per node and the available computational resources become diluted, leading to a gradual saturation of performance rather than degradation. Even as the network scales up, the proposed scheme maintains stable utility and balanced resource consumption under the given budget constraints. The results suggest that our scheme consistently outperforms other strategies across all node numbers, indicating an optimized balance between cost and quality incentives to maximize utility within the given budget. The quality-first and price-first strategies show competitive performance, although neither dominates across all node numbers. The randomized pricing and randomized node combination strategies yield lower utility, which may suggest a less effective node selection or pricing model under these approaches.
@@
(a) $\tau=9,\lambda=1$
(b) $\tau=5,\lambda=5$
(c) $\tau=1,\lambda=9$
Figure 6: Scheme performance under different preferences.
@@
Figure 8: Normalized server strategies for different parameters and unit cost of node.
Fig. 6 illustrates the performance comparison of different schemes under different preferences. Fig. 6(a) illustrates a situation where the user’s preference for quality far outweighs the concern for latency. With this preference, the satisfactionfirst strategy minimizes the value of AoI, indicating that the information is updated very frequently and that it is likely to perform well in providing quality services. At the same time, the price first strategy maintains the smallest data size, which indicates that the economy of data transmission is achieved even when quality is the primary concern. Fig. 6(b) illustrates the case where the user’s preference for both quality and latency is not strong. In this setting, the strategies are close in terms of AoI and latency, but the price first strategy performs the smallest in terms of data size, which is consistent with the observation in Fig. 6(a). Fig. 6(c) shows the setting where users prefer low latency. The price first strategy performs minimally in terms of data size, which may imply that fewer data are transmitted in the price first strategy in order to reduce the cost. The quality first strategy has a lower AoI, indicating that it provides more frequent information updates.
Fig. 7 illustrates the comparison of normalized server utility under different budget constraints. The figure shows three bar sets corresponding to different maximum budget constraints: $R^{max}=50$, $R^{max}=100$, and $R^{max}=150$. The number of selected nodes ranges from 5 to 25. As the number of selected nodes increases, the server utility generally increases under each budget constraint. However, the extent of the utility increase varies with the budget limit. For the lowest budget ($R^{max}=50$), the utility increases with the number of nodes but then levels off, indicating that beyond a certain point, additional nodes do not increase utility within the budget constraint. As the budget constraint increases from $R^{max}=100$ to $R^{max}=150$, utility continues to increase with the number of nodes, suggesting that higher budgets can effectively utilize more nodes to generate utility. The highest utility is observed under the constraint $R^{max}=150$, particularly with 25 nodes. This means that the server can utilize a higher budget to maximize its utility potential. The server’s ability to generate utility is significantly affected by the number of nodes it can support, which in turn is limited by its budget.
Fig. 8 illustrates the comparative analysis of server strategies under different conditions. The unit cost per node ranges from 1 to 5. Fig. 8 shows the server strategies under different unit satisfaction margins $\sigma$. As the node unit cost increases, the server strategies decrease for all values of $\sigma$. Higher satisfaction margins result in a sharp decrease in bids. This suggests that as the node cost increases, servers are inclined to bid higher for nodes with higher satisfaction margins. Fig. 8 shows the server’s bid for different data sizes $d$. As in Fig. 8, all three lines exhibit a decreasing trend, indicating that as node cost increases, the server’s willingness to bid declines. This decrease is more significant at larger data volumes, implying that servers favor nodes with larger data sizes when bidding. Fig. 8 shows that both unit satisfaction profit and data size significantly affect server strategy. As the node cost increases, servers tend to place lower bids, and higher satisfaction and larger data sizes amplify this effect.
Figure 9: FL training time for different number of selected nodes.
Figure 10: FL accuracy for different number of selected nodes.
@@ VII-C Solution Comparison
Fig. 11 compares the performance of different DRL algorithms in the Stackelberg game. In Fig. 11, as the number of selected clients increases, the server utility exhibits a rise-then-fall trend, reaching its peak at 15 clients. This trend occurs because a moderate number of clients provides sufficient computational resources to optimize training performance. However, as the number of clients continues to increase, the limited total budget results in less resource allocation per node, leading to a decline in overall utility. Of the four algorithms, MADDPG consistently achieves the highest utility, closely approximating the SE. In contrast, the utilities of MAPPO, MASAC, and MADQN are slightly lower, showing a noticeable gap. This demonstrates that MADDPG excels in balancing resource allocation and client selection to maximize server utility. In Fig. 11, as the node unit cost increases, server strategies gradually decrease because the server must lower its bids to optimize resource allocation under budget constraints. The server strategies in MADDPG remain consistent with the SE, whereas the server strategies in MAPPO, MASAC, and MADQN diverge from the SE. Fig. 11 shows that MADDPG consistently approximates the SE more closely than MAPPO, MASAC, and MADQN, indicating its superior performance in this Stackelberg game setting.
@@ VIII Conclusion
In this paper, we have designed the MEFL framework for IIoT. The framework leverages meta-computing to enhance FL in IIoT. We have developed a satisfaction-aware incentive scheme within MEFL. This scheme integrates a satisfaction function that incorporates data size, AoI, and latency into the utility function, and formulates the utility optimization problem as a two-stage Stackelberg game. We employ a DRL-based approach to learn the equilibrium of the Stackelberg game, aiming to maximize the server’s utility by identifying the optimal equilibrium. Experimental results demonstrate that, compared to traditional methods, the scheme effectively balances model quality and update latency through efficient resource allocation, enhances FL performance, ensures real-time capability and high efficiency of FL in IIoT, and better addresses the specific needs of industrial scenarios. However, the proposed MEFL still assumes rational and reliable participant behavior. Fully adversarial scenarios, such as intentional falsification of updates or coordinated manipulations, remain beyond the current scope. In future work, we plan to extend the framework by incorporating trust evaluation and anomaly detection mechanisms to further enhance its reliability and security in real-world IIoT environments.
[Uncaptioned image] Xiaohuan Li (Member, IEEE) was born in Chongqing, China. He received the B.Eng. and M.Sc. degrees from the Guilin University of Electronic Technology, Guilin, China, in 2006 and 2009, respectively, and the Ph.D. degree from the South China University of Technology, Guangzhou, China, in 2015. He was a Visiting Scholar with the Université de Nantes, France, in 2014. He is currently a Professor with the School of Information and Communication, Guilin University of Electronic Technology and Research fellow with the National Engineering Laboratory of Application Technology of Integrated Transportation Big Data (Beihang University). He is on the Editorial Boards of IEEE Internet of Things Journal and Journal on Communications. His current research interests include wireless sensor networks, vehicular networks, and cognitive radios.
[Uncaptioned image] Shaowen Qin received the B.S. degree from Guilin University of Electronic Technology, Guilin, China, in 2018, and the M.S. degree from Guangxi University, Nanning, China, in 2024. She is currently pursuing the Ph.D. degree in information and communication engineering at Guilin University of Electronic Technology. Her main research interests include federated learning and the industrial internet of things.
[Uncaptioned image] Xin Tang received the B.S. and M.S. degrees in 2011 and 2015, respectively, from the Guilin University of Electronic Technology, Guilin, China, where he is currently working toward the Ph.D. degree in information and communication engineering. Since 2015, he has been with the China Mobile Communications Corporation Guangxi Branch. In 2016, he joined the Institute of Information Technology, Guilin University of Electronic Technology, where he is currently a Senior Engineer. He was engaged in a graduate study abroad program at Nanyang Technological University, Singapore, from July 2024 to July 2025. His research interests include edge computing, multi-agent systems, UAV networks, and intelligent transportation systems.
[Uncaptioned image] Jiawen Kang received the Ph.D. degree from Guangdong University of Technology, China, in 2018. He has been a postdoc at Nanyang Technological University, Singapore from 2018 to 2021. He currently is a full professor at Guangdong University of Technology, China. His research interests mainly focus on blockchain, security, and privacy protection in wireless communications and networking.
[Uncaptioned image] Jin Ye received the Ph.D. degree with School of Science and Engineering, Central South University in 2008. She is currently a professor with school of Computer, Electronics and Information, Guangxi University. She worked as a visiting scholar in Department of Computer Science and Engineering at the University of Minnesota, Twin Cities in 2018. She also is the Member of China Computer Federation. Her main research interests include network protocol design, data center networks.
[Uncaptioned image] Zhonghua Zhao received the M.S. degree from Guilin University of Electronic Technology, Guilin, China. He is currently a professor with the School of Information and Communication, Guilin University of Electronic Technology, Guilin, China, and has published 18 academic papers. His current research interests include vehicular networks, edge computing, and intelligent transportation systems.
[Uncaptioned image] Yusi Zheng received the B.S. degree from Guilin University of Electronic Technology, Guilin, China, in 2024, and the M.S. degree from Lingnan University, Hong Kong, China, in 2025. He is currently pursuing the Ph.D. degree in information and communication engineering with the Guilin University of Electronic Technology. His research focuses on intelligent transportation systems, with applications in highway traffic event detection and UAV-based solutions.
[Uncaptioned image] Dusit Niyato (Fellow, IEEE) is a professor in the School of Computer Science and Engineering, at Nanyang Technological University, Singapore. He received B.Eng. from King Mongkuts Institute of Technology Ladkrabang (KMITL), Thailand in 1999 and Ph.D. in Electrical and Computer Engineering from the University of Manitoba, Canada in 2008. His research interests are in the areas of sustainability, edge intelligence, decentralized machine learning, and incentive mechanism design.